Dublin University Rifle Club

Version Date : May 21, 2001 Filename : Status :

DURC Match Secretary

Budget 2001-2002

Mark Dennehy May 21, 2001 Contents

Executive Summary 1

1 Introduction 2

2 Internal Competitions 2

2.1 Championships . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 2.1.1 Tournament Championships . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 2.1.2 Shoulder-to-Shoulder Championships . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2

2.2 Ladders . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 2.3 RO Knockout . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 2.4 Novelty Competitions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

2.4.1 Turkey Shoot . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 2.4.2 Rapid Fire Competition . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 2.4.3 Fencing Grudge Match . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

2.5 Three-Position Air Rifle . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

3 External Competitions 3

3.1 Postal Competitions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 3.1.1 Trinity Top Ten . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 3.1.2 Hartford Rematch . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4

3.2 Open Competitions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 3.2.1 DURC 10m Air Rifle Open . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 3.2.2 DURC 50m Prone Rifle Open . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4

3.3 RRPC vs DURC Competition . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 3.4 Colours Match . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 3.5 WTSC/UCDRC/DURC Competition . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4

List of Figures Executive Summary

1 2 Internal Competitions

1 Introduction

Given the experience from the previous two years I have prepared the following budget proposal for the DURC year from May 2001 to October 2002. This document should be reviewed and kept on file by the Treasurer of the club. Similar documents should be submitted by the Ammo&Equipment officer and other committee members and the Treasurer should draw up a preliminary budget based on these documents.

A note for this proposal is that we are in the process of applying for corporate sponsorship: Eircom have already been contacted and should they not provide the level of finance desired, other sponsors should be sought out. The prizes and expenditure levels in this proposal assume that funding from sponsors on the order of $5,000 has been secured.

Further, the proposed new classification scheme and the increased number of junior competitiors in Air Rifle has led to a large increase in the number of prizes required. Measures to economically provide worthwhile prizes have been discussed before. Medals and trophies can be fitted with custom centers with the DURC logo. These however, have to be obtained in a large batch: however, 1000 centers will last the club for a very long time, so that writing the cost off over several years of savings in prizes is a valid act.

Finally, this year is the 40^{th} anniversary of DURC’s founding: competitions thus take on a higher level of importance than in previous years.

2 Internal Competitions

The club has not run the full number of internal competitions in some years due to the slump of the past four years. Last year saw the membership drive and RO drive succeed in reversing this slump: now we must strive to restore the number of competitive events to its previous level.

2.1 Championships

Championships have been run in two different formats over the years : one-day and tournament style. I wish to run both, but modifing the one-day format to be a one-week format to allow for the maximum attendance possible.

2.1.1 Tournament Championships

These will be run in Hillary Term. We select the top 32 people from the ladders and fill any remaining places by invitation. Matches are made on the basis of the averages of the past ten cards (or as many are as available). Two weeks are allocated for two ladder cards to be shot. The highest gunscore wins and progresses to the next round. Results are to be placed on the noticeboard as they become available. The final will be shoulder-to-shoulder on a night both competitors agree on in the final week of term.

These rules apply to Air, Free and Sporter: though it is not expected that Sporter will attract as many entrants. This is not meant to be a disencouragement to the competition however.

Medals to be awarded to the top four in the competition; A perpetual shield should be obtained for the overall winner, but it should be displayed in the Club and a small tray given to the shooter.

2.1.2 Shoulder-to-Shoulder Championships

These will be run in Trinity Term. Open to any club member who wishes to take part. One week to be set aside for shooting Air Rifle, and one for Prone Rifle and Sporter. This will allow for Club Members to take part on the night most convienent for them. Reentrys should be considered; the tradeoff between extra entries versus percieved fairness must be made. Results should again be posted as they become available. Since an entire week is available, two ladder cards are to be shot for the championship.

Medals to be awarded to the top three in each discipline.

2.2 Ladders

2.2 Ladders

The ladders are central to the clubs operation: they act as a visual indication of how well shooters are progressing; they gauge the level of activity in the club; they serve as a long-running means of competition within the club; and they serve as a metric for evaluating the overall level of shooting standard in the club at year’s end.

Hence, the prizes should be more substantial than medals.

Discussion has taken place on the format of the ladders: more discussion is required at committee meetings to finalise the format.

2.3 RO Knockout

This competition was proposed last year but not implemented due to workload problems on committee. Last years comments still stand as valid:

This is a new competition to try to encourage RO currency in all disciplines. It will take place over Hilary term, and all disciplines will be shot. Entry is mandatory for all Range Officers. One card to be shot per week, in a nominated discipline at any point in the week. The prize to be awarded at the AGM.

2.4 Novelty Competitions

These competitions are not by any means a standard ISSF competition, but they do provide opportunity for excellent social occasions, which is important for the club’s membership.

2.4.1 Turkey Shoot

Prizes here are seen as a pleasant bonus - the emphasis is on the social function run concurrently with the shoot. Hence, simple medals for the winner and runner-up will suffice, though a seperate novice section should be considered.

2.4.2 Rapid Fire Competition

The comments in 2.4.1 apply equally here.

2.4.3 Fencing Grudge Match

Though not held last year due to workload on committee, this is a match with historical precendent dating back many years and should be held next year. Prizes are not usually given. Commerative medals should however be considered, and hospitality funds reserved for the post-match social function.

2.5 Three-Position Air Rifle

This competition has not been shot in several years, yet the perpetual shield remains in our possession. We should run this competition again this year, with medals for the winner and runner-up. A single night of shooting (or perhaps two) should prove sufficent.

3 External Competitions

Due to the special nature of this year, it is to be expected that a rise in the number of competitions will be seen.

3 External Competitions

3.1 Postal Competitions

UCESSA 25yd Postal

50m Postal 10m Air Postal

NSRA/Eley NSRA postal
TTT See 3.1.1
Hartford Rematch See 3.1.2

Several competitions are planned this year. Costs will be based on the entry fees, along with postal charges: all cards this year to be sent via registered mail. Note that this year sanctions will be in place for any shooter who disqualifies him or herself due to not meeting deadlines; that shooter will be obliged to repay their entry fee to the club.

3.1.1 Trinity Top Ten

As the Top Ten is regarded by the NTSA as the national postal, it must be run efficently. As such, this years funds must not be released for general Club use until such time as the prizes for the Top Ten have been paid for from them. Last years figures indicate that the Top Ten is a self-financing competition with healthy levels of interest: it had a small surplus of income last year. This year sees the introduction on a trial basis of a threeleague competition (10m Air Rifle, 25yd Prone and 50m Prone), with a Juniors section. This is to promote the level of entry to the competition and secure it’s status as the national postal. We may need to recruit scorers from the ROs if the workload proves too much for the Match Secretary alone. Should this happen, a maximum of two ROs will be assigned to one league, with one managing scoring of senior shooters and one of junior shooters; these ROs must accept this responsibility for the duration of the competition. The Match Secretarty will be responsible for coordination and recording in this case.

3.1.2 Hartford Rematch

In 1967, we won a competition via postal entries against Trinity College Hartford. They still have an active Rifle club and have been personally invited to enter this year’s TTT: however, I wish to run a rematch of the original postal competition this year as a commerative event. TCH were our first american opponents and the history says we won; this is hoped to be the impetus to get them to agree to a rematch.

A suitable engraved plate should be used as the prize for the winning team; commemorative medals should be issued to both our team and theirs. Ideally, costs for this should be divided between both colleges.

3.2 Open Competitions

3.2.1 DURC 10m Air Rifle Open

3.2.2 DURC 50m Prone Rifle Open

3.3 RRPC vs DURC Competition

3.4 Colours Match

3.5 WTSC/UCDRC/DURC Competition